One hundred students were asked about two kinds of music; jazz and classical. Sixty students said they listed to jazz while 45 said they listened to classical, and 20 didn't listen to either. How many students llistened to both styles?: One hundred students were asked about two kinds of music; jazz and classical. Sixty students said they listed to jazz while 45 said they listened to classical, and 20 didn't listen to either. How many students llistened to both styles?
Answer by vleith(1174)

Twenty don't listen at all. So that measn 100 - 20 = 80 do listen
60 are jazz, 80-60 = 20 who are classical only.
45-20=25 left must listen to both